OpenStudy (decarr432):

How do I find the radical form of this expression I will write it how it is Please explain slowly will fan and metal

OpenStudy (decarr432):

\[(32a ^{10}b ^{5 \over 2})^{2 \over 5}\]

OpenStudy (decarr432):

Anyone know how to solve?

zepdrix (zepdrix):

When you have a "fraction" or "rational" exponent, example: \(\large\rm x^{m/n}\) The numerator of this fraction represents the "power" as you're used to, while the denominator represents the degree of the "root" so in that example, I could rewrite it in this form: \(\large\rm \sqrt[n]{x^m}\)

OpenStudy (decarr432):

Okay

zepdrix (zepdrix):

So in your problem, we're focusing on the outermost exponent. The numerator is a 2, the "power" on the expression. while the denominator is a 5, the degree of our "root".

OpenStudy (decarr432):

Okay so it would make the answer A. but please continue

OpenStudy (decarr432):

i'd like to learn

zepdrix (zepdrix):

Good, option A :) \(\large\rm (stuff)^{2/5}=\sqrt[5]{(stuff)^2}\)

OpenStudy (decarr432):

Okay because the fraction outside the parenthesis would make up the radical?

zepdrix (zepdrix):

Outside the parenthesis? Well, only the denominator of that fraction is giving us the radical :) It becomes a little more tricky when they give you a 1 in the numerator. Realize that that represents a 1 "power", which is insignificant, so it really only represents a root. Example: \(\large\rm (stuff)^{1/2}=\sqrt[2]{(stuff)^1}=\sqrt{stuff}\) We'll relate this to the last problem in just a sec.

OpenStudy (decarr432):

Okay

zepdrix (zepdrix):

I'm not sure if you've learned all of your exponent rules at this point, but here is one of them,\[\large\rm (a^b)^c=a^{bc}\]In your original problem, we can actually apply this rule in reverse,\[\large\rm (stuff)^{2/5}=(stuff)^{2\cdot\frac{1}{5}}=\left[(stuff)^2\right]^{1/5}=\sqrt[5]{(stuff)^2}\]If that's way too complicated, you can ignore it for now :) Just another way to understand where this weird change from fraction to root is coming from.

OpenStudy (decarr432):

Okay ill keep that in mind actually

zepdrix (zepdrix):

Hmm I'm not sure what else we could say about this XD

OpenStudy (decarr432):

Thats fine im stuck on a new problem lol